OpenStudy (anonymous):

Questions & Anwsers!? 1...6x/7-x/2 =5 x = 10 x = 5 x = 14 2...3/x-1 + 2/x =4 {0.25, 2} {1/2, 2} {1/2, 2}

jimthompson5910 (jim_thompson5910):

# 1 6x/7-x/2 =5 14*6x/7-14*x/2 = 14*5 2*6x-7*x = 70 12x-7x = 70 5x = 70 x = 70/5 x = 14

jimthompson5910 (jim_thompson5910):

Basic idea: multiply everything by the LCD to clear out the fractions. Then solve like normal

jimthompson5910 (jim_thompson5910):

Use the same idea to solve #2

jimthompson5910 (jim_thompson5910):

Let me know how it goes

OpenStudy (anonymous):

i got 1

jimthompson5910 (jim_thompson5910):

Is the problem in #2 \[\Large \frac{3}{x-1}+\frac{2}{x} = 4\] ???

OpenStudy (anonymous):

ya

jimthompson5910 (jim_thompson5910):

This one is trickier than I thought, one moment

OpenStudy (anonymous):

ok i got 1/2 and 2

jimthompson5910 (jim_thompson5910):

I'll show you how to do this one

OpenStudy (anonymous):

Ok

jimthompson5910 (jim_thompson5910):

\[\Large \frac{3}{x-1}+\frac{2}{x} = 4\] \[\Large 3x+2(x-1) = 4x(x-1)\] \[\Large 3x+2x-2 = 4x^2-4x\] \[\Large 5x-2 = 4x^2-4x\] \[\Large 0 = 4x^2-4x-5x+2\] \[\Large 0 = 4x^2-9x+2\] \[\Large 4x^2-9x+2 = 0\] \[\Large (x-2)(4x-1) = 0\] \[\Large x-2=0 \ \text{ or } 4x-1=0\] \[\Large x=2 \ \text{ or } x=\frac{1}{4}\] So the solution set is {0.25, 2} since 1/4 = 0.25